6. A TV dish antenna receives EMM waves of frequency 2.38 GHz trom a satellite 36000 km away b) What is the distamce between successive wave crests of the electric eld vector of the waves? Hew much time in milliv cond, does the sigual requrs to trarel tem the satellite to the deh antenn? me in milliseconds do
Answer #1

Frequency of EM wave (f) = 2.55GHz = 2.55×109Hz

(b) for EM wave, wavelength is = c/f = (3×108)/(2.55×109)

= 0.118meter

(a) distance (d) = 3600km = 3.6×106m

Velocity(v) =c = 3×108m/s

time (t) = d/c = 3.6×106/(3×108) = 12 msec